# Dispositions and Roles of Generically Dependent Entities

### v1

@@ 4 Proposed changes to BFO
This section presents a possible approach that would enable BFO to support roles and dispositions of generally dependent continuants. Since the BFO developers tend to be conservative and, thus, like to avoid radical changes, the proposal is intended to minimize the necessary change. Hence, the following definitions and elucidations are based on BFO 2020 and only propose minor tweaks to most (if applicable, highlighted in bold font).
The following definition omits the domain and range restriction from BFO 2020.
@@
$x$ specifically depends on $y$ if and only if (a) $x$ is of a nature that, at any time $t$, $x$ cannot exist at $t$ if $y$ does not exist at $t$, (b) $x$ and $y$ do not share any parts and (c) $x$ is not a boundary of $y$.
The definition of inherence is modified by including generally dependent continuants.
@@
x inheres in y if and only if (a) x is a specifically dependent continuant and (b) y is an independent continuant that is not a spatial region or y is a generally dependent continuant and (c) x specifically depends on y.
The elucidation of realizable entity in BFO 2020 asserts that they are particulars of a type that is instantiated by some instances, which are realized in a corresponding process. The elucidation also asserts that these particulars inhere in independent continuants. This may be generalized to generally dependent continuants.
@@
A realizable entity is a specifically dependent continuant that (a) inheres in some independent continuant (which is not a spatial region) or inheres in a generally dependent continuant and that (b) is of a type some instances of which are realized in processes of a correlated type.
The realization of a realizable entity of a generally dependent continuants necessarily involves a concretization of the generally dependent continuant, either as part of the process or as a dependent continuant that inheres in a carrier of the generally dependent continuant. E.g., the disposition of stories to be scary is realized in processes that involve frightened listeners. Each of these processes involves a concretization of the scary story (e.g., by being printed in book or by telling the story).
@@
If a process $p$ is a realization of a realizable entity that inheres in a generally dependent continuant $x$, then either (a) a concretization of $x$ is part of $p$ or (b) some participant of $p$ is a carrier of $x$.
The notion of disposition and role needs to be generalized in BFO in order to support entities without a physical make-up. For this purpose, we also consider the retainers of sites as well as their parts and their arrangement. For example, the disposition of a cave to collapse may only cease to exist if its retainer (i.e., the surrounding rock) changes. Further, the disposition of a story to scare may only change if some of its parts are changed, e.g., by rewriting certain sections.
@@ 5 Conclusion
In this paper, we discussed some of the issues related to dispositions and roles of generally dependent continuants in BFO 2020. We discussed two approaches to solving them. The first, as proposed by Barry Smith, is the introduction of defined classes. As we have seen, this approach has serious limitations. Further, we suggested a few minor changes to BFO, which would address the issue directly and introduce dispositions and roles of generally dependent continuants.

### v2

@@ 4 Approach B: Changing BFO
This section presents a possible approach that would enable BFO to support roles and dispositions of generically dependent continuants. Since the BFO developers tend to be conservative and, thus, like to avoid radical changes, the proposal is intended to minimize the necessary change. Hence, the following definitions and elucidations are based on BFO 2020 and only propose minor tweaks (highlighted in bold font).
The definitions of specific dependence and inherence are modified by including generically dependent continuants in the range of the relationships.
@@
$x$ specifically depends on $y$ if and only if (a) $x$ is a specifically dependent continuant, (b) $y$ is a specifically dependent continuant or a generically dependent continuant or an independent continuant that is not a spatial region, (c) $x$ is of a nature that, at any time $t$, $x$ cannot exist at $t$ if $y$ does not exist at $t$, (d) $x$ and $y$ do not share any parts and (e) $x$ is not a boundary of $y$.
@@
x inheres in y if and only if (a) x is a specifically dependent continuant and (b) y is a generically dependent continuant or y is an independent continuant that is not a spatial region and (c) x specifically depends on y.
The elucidation of realizable entities in BFO 2020 asserts that they are particulars of a type that is instantiated by some instances, which are realized in a corresponding process. The elucidation also asserts that these particulars inhere in independent continuants. This may be generalized to generically dependent continuants.
@@
A realizable entity is a specifically dependent continuant that (a) inheres in some independent continuant (which is not a spatial region) or inheres in a generically dependent continuant and that (b) is of a type some instances of which are realized in processes of a correlated type.
The realization of a realizable entity of a generically dependent continuants necessarily involves a concretization of the generically dependent continuant, either as part of the process or as a dependent continuant that inheres in a carrier of the generically dependent continuant. E.g., the disposition of stories to be scary is realized in processes that involve frightened listeners. Each of these processes involves a concretization of the scary story (e.g., by being printed in a book or by telling the story).
@@
If a process $p$ is a realization of a realizable entity that inheres in a generically dependent continuant $x$, then either (a) a concretization of $x$ is part of $p$ or (b) some participant of $p$ is a carrier of $x$.
The notions of disposition and role need to be generalized in BFO in order to support entities without a physical make-up. For this purpose, we also consider parts and their arrangement, as well as retainers of immaterial entities. A retainer of site is a material entity in relation to which the site is defined ([6], pp. 112-113). For example, your office is a site, and its retainer is the aggregation of its floor, walls, and ceiling. In [6] the term ‘retainer’ is only introduced in the context of sites. However, since a continuant fiat boundary is also a kind of immaterial entity whose location is determined in relation to some material entity, we suggest to generalize its use for both sites and boundaries: a material object $x$ is a retainer of a site or a continuant fiat boundary $y$ if the the location of $y$ is determined in relation to $x$.
Considering the mereotopology and the retainers of entities enables us to generalize the definition of Role and Disposition in a way that covers sites, and continuant fiat boundaries, and generically dependent continuants.
@@
For example, the disposition of a cave to collapse may only cease to exist if its retainer (i.e., the surrounding rock) changes. However, the Geographic North Pole serves as a reference point for navigation not by disposition but by role, since it may lose this role without any alteration to its retainer (i.e., Earth). Further, the disposition of ‘Hansel and Gretel’ to scare children may only change if some of its parts are changed, e.g., by rewriting certain sections. However, its role as a running example of this paper was gained and may cease to exist without changing the mereotopology of the fairy tale.
@@ 5 Conclusion
In this paper, we discussed some of the issues related to dispositions and roles of generically dependent continuants in BFO 2020. We discussed two approaches to solving them. The first, as proposed by Barry Smith, is the introduction of defined classes. As we have seen, this approach has serious limitations. Further, we suggested a few minor changes to BFO, which would address the issue directly and introduce dispositions and roles of generically dependent continuants. The suggested changes have the additional benefit of addressing existing problems of BFO 2020 with respect to roles and dispositions of sites and continuant fiat boundaries.
Acknowledgements.
I would like to thank Barry Smith, Alan Ruttenberg, and the anonymous reviewers for their insightful comments. This work was funded by the Federal Ministry of Education and Research (BMBF) – funding code 13XP5187E.
Declaration on Generative AI
During the preparation of this work, the author used Grammarly in order to: Grammar and spelling check, paraphrase and reword. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the publication’s content.